OpenStudy (lucaz):

(2cm)^2 is equal to 4cm or 4cm^2 ?

zepdrix (zepdrix):

\[\Large (3x)^2\quad=\quad 3^2x^2\quad=\quad 9x^2\]We square everything in the brackets, even the units.\[\Large (2cm)^2\quad=\quad 2^2cm^2\]

OpenStudy (jdoe0001):

\(\bf (ab)^2\implies a^2b^2\)

OpenStudy (lucaz):

thanks